# Natural Language Can Help Bridge the Sim2Real Gap

### v1

@@ Abstract
The main challenge in learning image-conditioned robotic policies is acquiring a visual representation conducive to low-level control. Due to the high dimensionality of the image space, learning a good visual representation requires a considerable amount of visual data. However, when learning in the real world, data is expensive. Sim2Real is a promising paradigm for overcoming data scarcity in the real-world target domain by using a simulator to collect large amounts of cheap data closely related to the target task. However, it is difficult to transfer an image-conditioned policy from sim to real when the domains are very visually dissimilar. To bridge the sim2real visual gap, we propose using natural language descriptions of images as a unifying signal across domains that captures the underlying task-relevant semantics. Our key insight is that if two image observations from different domains are labeled with similar language, the policy should predict similar action distributions for both images. We demonstrate that training the image encoder to predict the language description or the distance between descriptions of a sim or real image serves as a useful, data-efficient pretraining step that helps learn a domain-invariant image representation. We can then use this image encoder as the backbone of an IL policy trained simultaneously on a large amount of simulated and a handful of real demonstrations. Our approach outperforms widely used prior sim2real methods and strong vision-language pretraining baselines like CLIP and R3M by 25 to 40%.
@@ I Introduction
A popular paradigm in foundation model research is to first pretrain the backbone on large datasets, and then add and train a task-specific head to process the backbone outputs to perform a downstream task. We borrow from this paradigm by first pretraining an image encoder to predict the pretrained embeddings of language descriptions of images from roughly $50$ trajectories in sim and real, with language labels on each image. Then we use this image encoder as the backbone of our IL policy and train on action-labeled data from both the sim and real domains simultaneously, where we only need a few action-labeled demonstrations from the real world.
@@ II Related Work
Our main contribution is a method to learn domain-invariant image representations by exploiting natural language descriptions as a bridge between domains for sim2real transfer. While we believe this has not been explored before, significant related research has been done in vision-language pretraining, domain-invariant representations for control, and sim2real techniques.
@@ II-A Vision Pretraining for Robotics
In vision-language pretraining, contrastive learning [41] has been shown to learn valuable representations for robotic tasks [50, 51]. However, these pretrained visual representations are often overly influenced by the semantics of language captions. This results in a representation that is too object-centric to differentiate between different frames of a robot demonstration, lacking the level of granularity needed for spatial-temporal understanding. R3M [36] addresses this by learning semantics from language labels of videos but also training with a time contrastive loss between video frames. Prior work in multimodal representations [63] found language to be effective in aligning representations learned across multiple modalities including depth and audio. Instead of using language to bridge modalities, our approach uses language to bridge visual representations between domains.
@@ II-B Sim2Real
While we approach sim2real through vision-language pretraining, there are many alternative, well-researched techniques. Domain randomization [4, 31, 54] involves varying physical parameters and visual appearances of the simulation to train a policy that functions in a wide distribution of domains that hopefully also covers the target domain distribution. However, domain randomization requires a large amount of diverse training data and attempts to be simultaneously performant in an overly broad distribution of states, leading to a suboptimal and conservative policy that takes longer to train. System identification [60, 26] involves tuning the simulation parameters to match the real world in order to create a custom-tailored simulation environment that easily transfers to the real domain. However, this process is very engineering intensive and time consuming, and it may be intractable to simulate all real world physical interactions with high fidelity and throughput. In contrast, our sim2real approach can handle potentially large discrepancies between the source and target domain and with a few target task demonstrations, does not require system identification or domain randomization.
@@ III Problem Description
All of our datasets are in the form of expert trajectories. Each trajectory, $\tau=\{(I_{t},s_{t},[a_{t},l_{t}],l_{task})\}$, is a sequence of tuples containing an image observation, $I_{t}$ ($128\times 128$ RGB), robot proprioceptive state, $s_{t}$ (end effector position and joint angles), and a language instruction of the task, $l_{task}$. Note that $l_{task}$ is the same over all timesteps of all trajectories in a given task. $[a_{t},l_{t}]$ denotes that a trajectory may optionally also include robot actions (in which case we consider the trajectory a full demonstration) and/or a language description of the image $I_{t}$. In the following sections, we identify with $\tau[L]$ a trajectory with language descriptions $l_{t}$, but no actions $a_{t}$. Similarly, $\tau[A]$ is a full demonstration with actions, $a_{t}$, but no language descriptions, $l_{t}$, and $\tau[AL]$ is a demonstration that contains both actions, $a_{t}$, and language, $l_{t}$.
@@ IV-A Automatic Language labeling of Images
To acquire image-language pairs for pretraining, we implement an automated pipeline for labeling the images of a trajectory that occurs synchronously during scripted policy demonstration collection (see Appendix VIII-A). Each if-case in the scripted policy corresponds to a stage index. We define a list of template strings describing the scene for each of these stages, so the stage indexes into the template string list, giving us our language annotation for the image. See Table III in the Appendix for all template strings, and Appendix VIII-G1 for details about our language labeling procedure.
However, our language labeling process need not be synchronously coupled with scripted policy demonstration collection. We also implemented a labeling process using off-the-shelf vision-language models to detect the location of objects and the gripper in the image to predict the stage number. This process can be run on previously-collected trajectories and requires only the images of a trajectory alone, without need for additional action or state information. We describe this second process in Appendix VIII-G2. Empirically, using language from this second, more scalable automated approach does not degrade the performance of our method.
@@ IV-B2 Language-Distance Learning
We also experiment with a second variant of image-language pretraining that incorporates language with a softer form of supervision. We posit that the exact values of the language embeddings do not themselves convey meaning; rather, it is the pairwise distances between two language embeddings that convey key information about the semantic similarity between two corresponding images. Thus, we design an objective to regress the image embedding distances between a pair of images from the two domains to their corresponding language distance:
@@
where the language distance function we use, $d:l\times l\rightarrow\mathbb{R}$ is BLEURT [46], a learned similarity score between two strings commonly used in the NLP community. We normalize $d(\cdot,\cdot)$ between 0 and 1 for all possible $\left(l^{s},l^{t}\right)$ pairs in our image-language dataset, where $1$ indicates the highest similarity between any two strings in the dataset. The output of $f_{cnn}$ is unit normalized before taking the dot product. When comparing both variants (see Sec. VI) we would like to assess if the additional degrees of freedom from the looser distance supervision are beneficial later on for policy training.
@@ V-A Sim2Sim and Sim2Real Environment Differences
In sim2sim, $\operatorname*{\operatorname*{\mathcal{D}}^{s}}$ and $\operatorname*{\operatorname*{\mathcal{D}}^{t}}$ are both sim environments with large differences in camera point-of-view (third person vs. first person), joint friction, and damping. In the sim2real setting, we employ a setup with a wide sim2real gap that we aim to bridge using language that includes differences in control frequency, task goals, visual observation appearance, objects, and initial positions. More details between the two environments in sim2sim and sim2real can be found in Appendix VIII-F.
@@ V-E Task Suite 2: Multi-step Pick and Place
In the sim2real setup, $\operatorname*{\operatorname*{\mathcal{D}}^{s}}$ remains the same, while $\operatorname*{\operatorname*{\mathcal{D}}^{t}}$ is the real task of putting a carrot into a bowl, then putting the bowl onto a plate (see Fig. 3), and ending with the gripper open after 50 timesteps. In addition to success rate (Section V-B), we measure average number of subtasks completed, allowing partial credit if the robot only succeeds in placing the carrot in the bowl, but not if only the bowl is placed on the plate.
@@ V-G Baselines
For the pretrained baselines, we consider two strong foundation models as the visual backbone, CLIP [41] and R3M [36], commonly used visual representations for robotics that are, like our approach, also shaped by language descriptions of images/videos.
@@ VI-A Experimental Questions and Analysis
What is the impact of our pretraining approach? Our method nearly doubles the success rate of both non-pretrained baselines in most task suites in sim2real and sim2sim. This indicates that Lang4Sim2Real can bridge a wide sim2real gap. One factor that may allow our method to perform well is that image observations with similar language descriptions may also have similar action labels. In Appendix VIII-C, we further investigate this hypothesis with an analysis of the action distributions between images, split by their language descriptions.
@@ VI-B Additional Experimental Questions and Results
What is the effect of pretraining on image-language pairs where the language granularity is reduced? We evaluate the impact of reduced language granularity on sim2real performance. See Appendix VIII-H for results.
How does our method perform if we cannot pretrain directly on image-language pairs from the target task? There are scenarios in which we might not have access to the real-world target task $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{target}}$ during the pretraining phase, as pretraining is often done without knowledge of the downstream task. To investigate this, we introduce a real-world prior task $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{prior}}$ that we pretrain on, and use real-world target task data $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{target}}$ only during imitation learning. The advantage of this problem setup is that we can reuse the same $f_{cnn}$ for multiple downstream real-world target tasks as long as they are sufficiently similar to the real-world prior task. In this modified problem setup, our method still mostly outperforms all baselines, which demonstrates that our method does not overfit to the real-world task it sees during pretraining. See Appendix VIII-I for full results.
@@ VIII Limitations and Future Work
One of the main limitations of our work is that the learned representation may have limited generalizability compared to existing pretraining methods that leverage internet-scale data to enable a large degree of generalization. Our approach targets a specific distribution and domain of real-world tasks and operates in the low-data regime for both pretraining and policy learning, so it does not yield general-purpose visual representations that can be applied to a wide distribution of target tasks.
@@
VIII-A Scripted Policy for Real-World Data Collection
@@
VIII-B Detailed Policy Network Architecture & Hyperparameters
@@
VIII-C Does Language Similarity Imply Action Distribution Similarity?
@@
VIII-D Task and Data Details
@@ VIII-F Sim2Sim and Sim2Real Differences
VIII-G Labeling Image Observations with Language
VIII-G1 Language labeling during Scripted Policy
@@
VIII-G2 Alternative Approach: Language labeling with off-the-shelf VLMs
To relax the requirement that our automated language labeling process must occur synchronously with a scripted policy collecting demonstrations, we implemented an alternative approach that is decoupled from the demonstration collection process. First, we use an off-the-shelf open-vocabulary object detector model, GroundingDINO [28], to output bounding boxes for the relevant objects on the scene. No finetuning of GroundingDINO is required. Second, we train a CNN-based gripper state predictor to predict the gripper position $(x,y,z)$ as well as whether the gripper is opened or closed in a given image. This network is trained on previously collected ((((image, gripper position, gripper opened/closed)))) data from 100 trajectories, and takes one minute to train on a single A5000 GPU. Using these two models, we can get the gripper state and position relative to the objects, enabling us to predict a stage number that corresponds fairly closely with the actual stage number as outputted by our scripted policy. Finally, we verified that training our method on VLM-derived language annotations does not degrade performance.
VIII-H Impact of Language Granularity on Performance
@@
VIII-I sim2real results with no pretraining on $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{target}}$

### v2

@@ Abstract
The main challenge in learning image-conditioned robotic policies is acquiring a visual representation conducive to low-level control. Due to the high dimensionality of the image space, learning a good visual representation requires a considerable amount of visual data. However, when learning in the real world, data is expensive. Sim2Real is a promising paradigm for overcoming data scarcity in the real-world target domain by using a simulator to collect large amounts of cheap data closely related to the target task. However, it is difficult to transfer an image-conditioned policy from sim to real when the domains are very visually dissimilar. To bridge the sim2real visual gap, we propose using natural language descriptions of images as a unifying signal across domains that captures the underlying task-relevant semantics. Our key insight is that if two image observations from different domains are labeled with similar language, the policy should predict similar action distributions for both images. We demonstrate that training the image encoder to predict the language description or the distance between descriptions of a sim or real image serves as a useful, data-efficient pretraining step that helps learn a domain-invariant image representation. We can then use this image encoder as the backbone of an IL policy trained simultaneously on a large amount of simulated and a handful of real demonstrations. Our approach outperforms widely used prior sim2real methods and strong vision-language pretraining baselines like CLIP and R3M by 25 to 40%. See additional videos and materials at https://robin-lab.cs.utexas.edu/lang4sim2real/.
@@ I Introduction
A popular paradigm in foundation model research is to first pretrain the backbone on large datasets, and then add and train a task-specific head to process the backbone outputs to perform a downstream task. We borrow from this paradigm by first pretraining an image encoder to predict the pretrained embeddings of language descriptions of images from roughly a few hundred trajectories in sim and real, with language labels on each image. Then we use this image encoder as the backbone of our IL policy and train on action-labeled data from both the sim and real domains simultaneously, where we only need a few action-labeled demonstrations from the real world.
@@ II Related Work
Our main contribution is a method to learn domain-invariant image representations by exploiting natural language descriptions as a bridge between domains for sim2real transfer. While we believe this has not been explored before, significant related research has been done in vision-language pretraining, sim2real techniques, and domain-invariant representations for control.
@@ II-A Vision Pretraining for Robotics
In vision-language pretraining, contrastive learning [42, 63] has been shown to learn valuable representations for robotic tasks [51, 52]. However, these pretrained visual representations are often overly influenced by the semantics of language captions. This results in a representation that is too object-centric to differentiate between different frames of a robot demonstration, lacking the level of granularity needed for spatial-temporal understanding. R3M [37] addresses this by learning semantics from language labels of videos but also training with a time contrastive loss between video frames. Prior work in multimodal representations [65] found language to be effective in aligning representations learned across multiple modalities including depth and audio. Instead of using language to bridge modalities, our approach uses language to bridge visual representations between domains.
@@ II-B Sim2Real
While we approach sim2real through vision-language pretraining, there are many alternative, well-researched techniques. Domain randomization [3, 32, 55] involves varying physical parameters and visual appearances of the simulation to train a policy that functions in a wide distribution of domains that hopefully also covers the target domain distribution. However, domain randomization requires a large amount of diverse training data and attempts to be simultaneously performant in an overly broad distribution of states, leading to a suboptimal and conservative policy that takes longer to train. System identification [61, 26] involves tuning the simulation parameters to match the real world in order to create a custom-tailored simulation environment that easily transfers to the real domain. However, this process is very engineering intensive and time consuming, and it may be intractable to simulate all real world physical interactions with high fidelity and throughput. In contrast, our sim2real approach can handle large source and target domain discrepancies with a few target task demonstrations and does not require system identification or domain randomization.
@@ III Problem Description
All of our datasets are in the form of expert trajectories. Each trajectory, $\tau=\{(I_{t},s_{t},[a_{t},l_{t}],l_{task})\}$, is a sequence of tuples containing an image observation, $I_{t}$ ($128\times 128$ RGB), robot proprioceptive state, $s_{t}$ (end effector position and joint angles), and a language instruction of the task, $l_{task}$. Note that $l_{task}$ is the same over all timesteps of all trajectories in a given task. $[a_{t},l_{t}]$ denotes that a trajectory may optionally also include robot actions (in which case we consider the trajectory a full demonstration) and/or a language description of the image $I_{t}$. In the following sections, we identify with $\tau[L]$ a trajectory with language descriptions $l_{t}$, but no actions $a_{t}$. Similarly, $\tau[A]$ is a full demonstration with actions, $a_{t}$, but no language descriptions, $l_{t}$.
@@ IV-A Automatic Language labeling of Images
To acquire image-language pairs for pretraining, we implement an automated pipeline for labeling the images of a trajectory that occurs synchronously during scripted policy demonstration collection (see Appendix IX-A). Each if-case in the scripted policy corresponds to a stage index, where in pick-and-place, the first stage corresponds to the gripper moving to a point above the object, the second stage corresponds to the gripper moving vertically down toward the object, and so on. We define a list of template strings describing the scene for each of these stages, so the stage indexes into the template string list, giving us our language annotation for the image. See Table III in the Appendix for all template strings, and Appendix IX-G1 for details about our language labeling procedure.
However, our language labeling process need not be synchronously coupled with scripted policy demonstration collection. We also implemented a labeling process using off-the-shelf vision-language models to detect the location of objects and the gripper in the image to predict the stage number. This process can be run on previously-collected trajectories and requires only the images of a trajectory alone, without need for additional action or state information. We describe this second process in Appendix IX-G2. Empirically, using language from this second, more scalable automated approach does not degrade the performance of our method.
@@ IV-B2 Language-Distance Learning
We also experiment with a second variant of image-language pretraining that incorporates language with a softer form of supervision. We posit that the exact values of the language embeddings do not themselves convey meaning; rather, key information about the semantic similarity of two images lies in the pairwise distances between their corresponding two language embeddings. Thus, we design an objective to regress the image embedding distances between a pair of images from the two domains to their corresponding language distance:
@@
where the language distance function we use, $d:l\times l\rightarrow\mathbb{R}$ is BLEURT [47], a learned similarity score between two strings commonly used in the NLP community. We normalize $d(\cdot,\cdot)$ between 0 and 1 for all possible $\left(l^{s},l^{t}\right)$ pairs in our image-language dataset, where $1$ indicates the highest similarity between any two strings in the dataset. Empirically, over our set of language descriptions, we found BLEURT provided a richer signal than simply taking dot products or $\ell$-2 distances between language embeddings. The output of $f_{cnn}$ is unit normalized before taking the dot product. We compare both variants (see Sec. VI) to assess whether the additional degrees of freedom from the looser distance supervision are beneficial later on for policy training.
@@ V-A Sim2Sim and Sim2Real Environment Differences
In sim2sim, $\operatorname*{\operatorname*{\mathcal{D}}^{s}}$ and $\operatorname*{\operatorname*{\mathcal{D}}^{t}}$ are both sim environments with large differences in camera point-of-view (third person vs. first person), joint friction, and damping. In the sim2real setting, we employ a setup with a wide sim2real gap that we aim to bridge using language that includes differences in control frequency, task goals, visual observation appearance, objects, and initial positions. More details between the two environments in sim2sim and sim2real can be found in Appendix IX-F.
@@ V-E Task Suite 2: Multi-step Pick and Place
In the sim2real setup, $\operatorname*{\operatorname*{\mathcal{D}}^{s}}$ remains the same, while $\operatorname*{\operatorname*{\mathcal{D}}^{t}}$ is the real task of putting a carrot into a bowl, then putting the bowl onto a plate (see Fig. 3), and ending with the gripper open after 50 timesteps. In addition to success rate (Section V-B), we measure average number of consecutive subtasks completed from the beginning, allowing partial credit if the robot only succeeds in the first step of placing the carrot in the bowl. However, if the robot does not finish the first step but finishes the second step, we do not count this as having completed any subtasks.
@@ V-G Baselines
For the pretrained baselines, we consider two strong foundation models as the visual backbone, CLIP [42] and R3M [37], commonly used visual representations for robotics that are, like our approach, also shaped by language descriptions of images/videos, and add a trainable policy head composed of fully-connected layers sharing the same dimensions as all other methods in our results.
@@ VI-A Experimental Questions and Analysis
What is the impact of our pretraining approach? Our method nearly doubles the success rate of both non-pretrained baselines in most task suites in sim2real and sim2sim. This indicates that Lang4Sim2Real can bridge a wide sim2real gap. One factor that may allow our method to perform well is that image observations with similar language descriptions may also have similar action labels. In Appendix IX-C, we further investigate this hypothesis with an analysis of the action distributions between images, split by their language descriptions.
@@ VI-B Additional Experimental Questions and Results
What is the effect of pretraining on image-language pairs where the language granularity is reduced? We evaluate the impact of reduced language granularity on sim2real performance. See Appendix IX-H for results.
How does our method perform if we cannot pretrain directly on image-language pairs from the target task? There are scenarios in which we might not have access to the real-world target task $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{target}}$ during the pretraining phase, as pretraining is often done without knowledge of the downstream task. To investigate this, we introduce a real-world prior task $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{prior}}$ that we pretrain on, and use real-world target task data $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{target}}$ only during imitation learning. The advantage of this problem setup is that we can reuse the same $f_{cnn}$ for multiple downstream real-world target tasks as long as they are sufficiently similar to the real-world prior task. In this modified problem setup, our method still mostly outperforms all baselines, which demonstrates that our method does not overfit to the real-world task it sees during pretraining. See Appendix IX-I for full results.
Can our method be combined with prior large-scale vision-language pretrained networks? We experiment with combining R3M and our method. Results are discussed in Appendix IX-J. Similar to [63], we find that finetuning R3M on our pretraining procedure does not yield improvements on downstream policy learning.
What are some failure modes of our policy? See our project website at https://robin-lab.cs.utexas.edu/lang4sim2real/ for failure videos. In stack object, most failures arose from not closing the gripper enough before attempting to lift the object, since our gripper action dimension is analog (controls the distance between the gripper fingers) instead of binary (open/closed). In multi-step pick-and-place, most failures of our policy came from the robot not moving right after successfully completing the first pick-and-place operation. In wrap wire, most failures were from grasping imprecision, since the small plug allowed for only a $\leq 3$ cm $xy$-plane error in the grasping point. Other common failures included prematurely opening the gripper and dropping the plug before the wrapping was finished, or moving the gripper sideways into the blender instead of properly circling around it.
@@ VIII Limitations and Future Work
One of the main limitations of our work is that the learned representation may have limited generalizability compared to existing pretraining methods that leverage internet-scale data to enable a large degree of generalization. Our approach targets a specific distribution and domain of real-world tasks and operates in the low-data regime for both pretraining and policy learning, so it does not yield general-purpose visual representations that can be applied to a wide distribution of target tasks. Future work could investigate scaling our method to large-scale datasets to further reduce the number of real-world demonstrations needed for effective sim2real transfer.
@@
IX Acknowledgements
We would like to thank members of the RobIn Lab at UT Austin for their valuable suggestions and help with debugging real robot issues. This research was supported by NSF NRI Grant IIS-1925082.
@@ IX Acknowledgements
IX-A Scripted Policy for Real-World Data Collection
@@
IX-B Detailed Policy Network Architecture & Hyperparameters
@@
IX-C Does Language Similarity Imply Action Distribution Similarity?
@@
IX-D Task and Data Details
@@ IX-F Sim2Sim and Sim2Real Differences
IX-G Labeling Image Observations with Language
IX-G1 Language labeling during Scripted Policy
@@
IX-G2 Alternative Approach: Language labeling with off-the-shelf VLMs
To relax the requirement that our automated language labeling process must occur synchronously with a scripted policy collecting demonstrations, we implemented an alternative approach that is decoupled from the demonstration collection process. First, we use an off-the-shelf open-vocabulary object detector model, GroundingDINO [29], to output bounding boxes for the relevant objects on the scene. No finetuning of GroundingDINO is required. Second, we train a CNN-based gripper state predictor to predict the gripper position $(x,y,z)$ as well as whether the gripper is opened or closed in a given image. This network is trained on previously collected ((((image, gripper position, gripper opened/closed)))) data from 100 trajectories, and takes one minute to train on a single A5000 GPU. Using these two models, we can get the gripper state and position relative to the objects, enabling us to predict a stage number that corresponds fairly closely with the actual stage number as outputted by our scripted policy. Finally, we verified that training our method on VLM-derived language annotations does not degrade performance. We performed image-language pretraining with language labels from either labeling method and tested on 2-step pick-and-place with 100 real-world trajectories. Both methods achieve 90% success rate averaged over 2 seeds.
IX-H Impact of Language Granularity on Performance
@@
IX-I sim2real results with no pretraining on $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{target}}$
@@
TABLE X: sim2real: Finetuning R3M with our method
Pretraining Action-labeled Data Multi-step Pick and Place Wrap Wire
Sim Real Success Rate (%) Subtasks Completed Success Rate (%)
$\operatorname*{\operatorname*{\mathcal{D}}^{s}}$ $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{target}}$ $\operatorname*{\operatorname*{\mathcal{D}}^{t}_{prior}}$ 25 50 100 25 50 100 25 50 100
R3M + adapters, Lang Reg. 0 30 40 0.7 0.95 1.25 0 5 5
Lang Reg. (ours) 55 80 95 1.2 1.8 1.95 25 50 55
R3M (frozen) 75 75 85 1.6 1.55 1.75 30 25 20
IX-J Combining R3M with Our Approach
We implemented and evaluated multiple ways to combine R3M with our image-language pretraining to see if it would be possible to leverage the benefits of both R3M’s large-scale pretraining and our method’s domain-invariant representation learning. Instead of initializing a ResNet from scratch before image-language pretraining, we experiment with using R3M weights and finetuning the last layer, the entire network, or inserted convolutional adapter modules [8]. Finetuning adapters (denoted R3M$+$adapters) performs the best in sim2sim, matching the performance of our method on 2-step pick-and-place.
Based on this sim2sim performance, we evaluated R3M$+$adapters on sim2real, but this performed worse than either frozen R3M or our method in sim2real (Table X). We hypothesize that this is because during image-language pretraining on both sim and real images, the trainable adapters learn to pick out features primarily in the simulation images as this is out-of-distribution for R3M which was trained on real-world videos, which enables R3M$+$adapters to do well in sim2sim but not sim2real.